# LabelCraft: Empowering Short Video Recommendations with Automated Label Crafting

### v1

@@ 3.3. Learning Strategy
The two levels of optimization in Equation (3.2) affect each other, forming a loop of nested optimization. Directly solving the total optimization problem is challenging. To overcome this challenge, we develop a training method based on the meta-learning technique — MAML (Finn et al., 2017), which considers alternatively updating the $\phi$ and $\theta$ in a loop. For each interaction, we update $\phi$ and $\theta$ as follows:
@@ 3.4. Platform Objectives
In this subsection, we outline our approach for evaluating a recommender model’s performance with respect to platform objectives. To address the expensive cost of the online experiments, we propose using a hold-out dataset $\mathcal{D}_{v}$ to achieve these goals, as represented by $M(f{\theta};\mathcal{D}_{v})$ in Equation (3a). As discussed in Section 1, the typical platform objectives consist of three sub-objectives: user usage time, engagement, and retention. We now describe how these sub-objectives can be computed using the soft top-$k$ technique and then describe the process of balancing them, considering their varying levels of learning difficulty.
@@ 4.4.1. Debiasing Performance
To assess the effectiveness of our method in mitigating duration bias, which is a notable strength shared by D2Q and DVR, we conduct an analysis experiment on Kuaishou following the approach outlined in the DVR paper (Zheng et al., 2022). Specifically, we analyze the histogram depicting the distribution of video duration in the top-$k$ recommended video lists generated by different models, and compare the results of the non-debiasing method PC, the debiasing methods D2Q and DVR, and our LabelCraft. The comparative results are summarized in Figure 2.
According to Figure 2, the PC method could not mitigate duration bias, as evidenced by its extremely skewed recommendation distribution and poor recommendation accuracy (see NWTG@10 and NEG@10 in Table 2). The recommendations of method DVR and D2Q become more balanced by debiasing as expected. Interestingly, our LabelCraft method demonstrates more balanced recommendations compared to DVR and D2Q. Additionally, it achieves better performance on accuracy-related metrics, as evidenced by the results on NWTG@10 and NEG@10 in Table 2. These findings indicate that LabelCraft effectively mitigates duration bias. The debiasing capability can be attributed to the alignment between the labels generated by LabelCraft and the platform objectives. When the platform objective is appropriately designed to be free of biases, it could guide the labeling model to generate bias-free labels.
@@ 4.4.2. The effect of Hyper-parameter $\tau$
In our dynamic balancing design described in Equation (12), the temperature coefficient ($\tau$) plays a crucial role in determining the strength of our dynamic balancing strategy. We conduct a study to investigate the impact of varying $\tau$ on LabelCraft’s performance. Figure 3 illustrates the performance of LabelCraft for each metric, along with the best performance achieved by baselines on each metric, as $\tau$ is varied within the range of $[0,1]$ with a step size of 0.1. From Figure 3, it is evident that when $\tau=0.5$, LabelCraft consistently outperforms all baselines across all evaluation metrics. In the range of $\tau\in[0.5,0.7]$, the performance of LabelCraft remains relatively acceptable, either surpassing or comparable to the best baseline results on each metric [^3] Within this range, LabelCraft achieves lower results on DS@10 compared to the best baseline. However, the impact on the balance of recommendation results remains small and the performance is close to the result of the second-best baseline.. However, it is noted that consistent superiority over the best baseline performance cannot be maintained within this range. This suggests the importance of setting an appropriate value for $\tau$.
(a) PC
(b) D2Q
(c) DVR
(d) LabelCraft
Figure 2. Distribution of video duration in the top-$k$ recommended video lists generated by the PC, D2Q, DVR, and our LabelCraft for Kuaishou.
@@ 5. Related Work
(a) NWTG@10 and NEG@10
(b) DS@10
Figure 3. Performance of LabelCraft across different values of $\tau$, compared to the best results achieved by baselines.

### v2

@@ 3.3. Learning Strategy
The two levels of optimization in Equation (3) affect each other, forming a loop of nested optimization. Directly solving the total optimization problem is challenging. To overcome this challenge, we develop a training method based on the meta-learning technique — MAML (Finn et al., 2017), which considers alternatively updating the $\phi$ and $\theta$ in a loop. For each interaction, we update $\phi$ and $\theta$ as follows:
@@ 3.4. Platform Objectives
In this subsection, we outline our approach for evaluating a recommender model’s performance with respect to platform objectives. To address the expensive cost of the online experiments, we propose using a hold-out dataset $\mathcal{D}_{v}$ to achieve these goals, as represented by $M(f{{}_{\theta}};\mathcal{D}_{v})$ in Equation (3a). As discussed in Section 1, the typical platform objectives consist of three sub-objectives: user usage time, engagement, and retention. We now describe how these sub-objectives can be computed using the soft top-$k$ technique and then describe the process of balancing them, considering their varying levels of learning difficulty.
@@ 4.4.1. Debiasing Performance
To assess the effectiveness of our method in mitigating duration bias, which is a notable strength shared by D2Q and DVR, we conduct an analysis experiment on Kuaishou following the approach outlined in the DVR paper (Zheng et al., 2022). Specifically, we analyze the histogram depicting the distribution of video duration in the top-$k$ recommended video lists generated by different models, and compare the results of the non-debiasing method PC, the debiasing methods D2Q and DVR, and our LabelCraft. The comparative results are summarized in Figure LABEL:fig:ana_duration.
According to Figure LABEL:fig:ana_duration, the PC method could not mitigate duration bias, as evidenced by its extremely skewed recommendation distribution and poor recommendation accuracy (see NWTG@10 and NEG@10 in Table 2). The recommendations of method DVR and D2Q become more balanced by debiasing as expected. Interestingly, our LabelCraft method demonstrates more balanced recommendations compared to DVR and D2Q. Additionally, it achieves better performance on accuracy-related metrics, as evidenced by the results on NWTG@10 and NEG@10 in Table 2. These findings indicate that LabelCraft effectively mitigates duration bias. The debiasing capability can be attributed to the alignment between the labels generated by LabelCraft and the platform objectives. When the platform objective is appropriately designed to be free of biases, it could guide the labeling model to generate bias-free labels.
@@ 4.4.2. The effect of Hyper-parameter $\tau$
In our dynamic balancing design described in Equation (12), the temperature coefficient ($\tau$) plays a crucial role in determining the strength of our dynamic balancing strategy. We conduct a study to investigate the impact of varying $\tau$ on LabelCraft’s performance. Figure LABEL:hyper_gamma illustrates the performance of LabelCraft for each metric, along with the best performance achieved by baselines on each metric, as $\tau$ is varied within the range of $[0,1]$ with a step size of 0.1. From Figure LABEL:hyper_gamma, it is evident that when $\tau=0.5$, LabelCraft consistently outperforms all baselines across all evaluation metrics. In the range of $\tau\in[0.5,0.7]$, the performance of LabelCraft remains relatively acceptable, either surpassing or comparable to the best baseline results on each metric [^3] Within this range, LabelCraft achieves lower results on DS@10 compared to the best baseline. However, the impact on the balance of recommendation results remains small and the performance is close to the result of the second-best baseline.. However, it is noted that consistent superiority over the best baseline performance cannot be maintained within this range. This suggests the importance of setting an appropriate value for $\tau$.